# Diffusion Graph Neural Networks for Robustness in Olfaction Sensors and Datasets in Security Robotics

### v3

@@ Abstract
Robotic odour source localization (OSL) is a critical capability for autonomous systems operating in complex environments. However, current OSL methods often suffer from ambiguities, particularly when robots misattribute odours to incorrect objects due to limitations in olfactory datasets and sensor resolutions. To address this challenge, we introduce a novel machine learning method using diffusion-based molecular generation to enhance odour localization accuracy that can be used by itself or with automated olfactory dataset construction pipelines. This generative process of our diffusion model expands the chemical space beyond the limitations of both current olfactory datasets and training methods, enabling the identification of potential odourant molecules not previously documented. The generated molecules can then be more accurately validated using advanced olfactory sensors, enabling them to detect more compounds and inform better hardware design. By integrating visual analysis, language processing, and molecular generation, our framework enhances the ability of olfaction-vision models on robots to accurately associate odours with their correct sources, thereby improving navigation and decision-making through better sensor selection for a target compound in critical applications such as explosives detection, narcotics screening, and search and rescue. Our methodology represents a foundational advancement in the field of artificial olfaction, offering a scalable solution to challenges posed by limited olfactory data and sensor ambiguities [^1] Code and data are made available to the community at the following URL: https://github.com/KordelFranceTech/OlfactionVisionLanguage-Dataset .
@@ I Introduction
The human sense of smell is a complex and nuanced sensory modality, capable of distinguishing an extensive array of odourants. In recent years, artificial olfactory systems have been developed to detect sources of explosives, illegal drugs, and even perform forensic studies. These olfactory systems, such as electronic noses, utilize sensor arrays and pattern recognition algorithms to detect and classify volatile compounds [1, 2, 3, 4].
Despite sensor advancements, the development of machine learning models for olfactory perception faces significant challenges, primarily due to the scarcity of comprehensive olfactory datasets. Humans describe chemical odourants with lingual descriptors such as fruity and floral, but sensors interpret odourants by their direct chemical reactions, leading to a disparity in data training methods for olfactory sensors (e.g. how does one describe to an olfactory sensor the scent of explosives?). Datasets like those from Leffingwell [5] and GoodScents [6] have been instrumental in providing foundational results. However, due to the complexity in obtaining human-labeled odour descriptors over compounds, these datasets are limited in scope, lacking complete coverage of the vast chemical space associated with olfactory stimuli [7] and the necessary human controls to ensure objective measurements such as compensation for genetic variation in olfactory mechanisms [8], environmental context [9], health status [10, 11], age [12], and even potential neurocognitive conditions [13].
To address this data paucity, we propose an innovative diffusion-based graph neural network that allows one to minimize uncertainty the construction of olfactory datasets, particularly olfaction-vision datasets. Our motivation is driven by the desire to teach robots to navigate by scent and triangulate the source of explosives, building on methods from [14, 15, 16, 17]. Our research provides four primary contributions:
@@
    Our model aids in the selection of olfactory sensors for a desired target compound.
@@
    In result of (1) and (2), our model reduces overall uncertainty and bias in olfactory navigation tasks.
@@
    As a result of (3), our methods directly increase accuracy and safety in olfaction-based security applications.
@@
II Background
The intersection of machine learning and olfaction presents a compelling frontier in artificial intelligence research. The human olfactory system’s ability to discern a vast array of scents is a complex process, involving intricate interactions between odourant molecules and olfactory receptors. Replicating this capability computationally necessitates sophisticated models and comprehensive datasets.
@@ II Background
A significant hurdle in developing machine learning models for olfaction is the scarcity of extensive and high-quality datasets. Existing resources, such as the Leffingwell PMP 2001 and GoodScents databases, provide valuable information on various chemical compounds and their associated odour descriptors. However, these datasets are limited in scope and do not encompass the full spectrum of olfactory stimuli. This limitation hampers the training of robust machine learning models capable of generalizing across diverse olfactory inputs.
@@
To address the data scarcity issue, researchers have explored the use of vision-language models (VLMs) to generate odour descriptors from images. VLMs, trained on large-scale image-text pairs, can infer contextual information from visual inputs and generate corresponding textual descriptions [18] [19]. By applying VLMs to images, it is possible to extract descriptive terms that may correlate with olfactory characteristics. However, the reliability of these descriptors is contingent upon the VLM’s exposure to olfactory-related data during training, the degree of which is not known. Consequently, the generated descriptors may lack specificity or accuracy in representing actual odours.
@@
While efforts to construct olfaction-vision-language models (OVLMs) exist [20, 21], their underlying training data and training processes are not well standardized. By analyzing visual content in a standard computer vision dataset like, for example, the COCO dataset [22], VLMs can infer potential olfactory characteristics associated with depicted scenes or objects. For example, if a motorcycle is identified in an image, the VLM will reason that carbon monoxide is likely a present molecule and will list the aromas that describe carbon monoxide given the appropriate prompt. These inferred descriptors serve as a bridge to map visual information to chemical compounds, facilitating the identification of molecules likely present in a given image.
@@
However, the reliance on VLMs introduces its own set of uncertainties, particularly regarding the accuracy and completeness of the generated odour descriptors. It is difficult to know the exact scope of chemistry contained in the VLM training data. To mitigate these concerns, we integrate a diffusion model trained on existing olfactory datasets and designed to generate novel molecular structures that correspond to the inferred odour descriptors, effectively expanding the chemical space beyond the limitations of current datasets. By doing so, we aim to capture a broader spectrum of odourant molecules, including those not previously documented. This is helpful in tasks oriented around olfactory sensors. Sensors that rapidly sample the air only screen for a select set of molecules. If a robot is given the task to navigate toward a specific odour, it has to be equipped with olfactory sensors that can target the specific molecules attributed to that odour. By understanding the target molecule’s near neighbors on the odour continuum, one can ensure to integrate sensors that address these molecules as well. As a simplified example, imagine that an unmanned aerial vehicle (UAV) is instructed to find the source of nitric oxide ($NO$), sometimes an indicator for the presence of explosives. $NO$ can, in some environments, quickly oxidize to $NO_{2}$. If the robot takes substantial time to navigate to the source of $NO$, it could be analyzing the wrong compound by missing the detection window, and at some point, may want to transition to detecting $NO_{2}$. The olfactory sensors (for example, metal oxide sensors) required to detect both $NO$ and $NO_{2}$ are different, and it would be prudent for the engineer to query which near neighbor compounds to $NO$ could be relevant for detection so the UAV could be equipped with such sensors. This issue is exemplified in tasks from [14] [15] [16] [17] [23]. We can make such a query automatic in our dataset construction, such that a separate machine learning model is queried for the nearest neighbors of any compound identified by the VLM. This not only informs better sensor selection, but can also partially reconcile erroneous compounds produced by the VLM.
@@
In machine olfaction, the Shape Theory of Olfaction suggests that molecules that exhibit similar structural traits ”smell” similarly [24] [25] [26] . Although there is some support against this theory [27] [28] [29] [30], our proposed method assumes this theory is true, and we construct a model that builds off its principles. For example, if a VLM indicates that the odour descriptors ”fruity” and ”floral” are associated with an object in an image, one can use the GoodScent or Leffingwell datasets to look up particular molecules that exude such aromas. However, since these datasets are not comprehensive of all possible molecules, one can use the proposed diffusion model to generate ”near neighbors” of molecules that are structurally similar to those found in the existing datasets.
Figure 1: An example of how diffusion helps in olfactory modeling. A VLM describes aromas present in COCO images. Near neighbors of those aromas and molecules are then synthesized with our diffusion model for more comprehensive olfactory analysis of the visual scene.
@@
To bridge the gap between odour descriptors and chemical compounds, diffusion models have emerged as a promising solution. Diffusion models are generative frameworks that learn to produce data samples by iteratively denoising random noise, guided by learned data distributions. In the context of molecular generation, diffusion models can be conditioned on textual inputs, such as odour descriptors, to generate novel molecular structures that correspond to the specified olfactory characteristics. This approach enables the exploration of chemical spaces beyond existing datasets, facilitating the discovery of new compounds with desired scent profiles. Recent studies have demonstrated the efficacy of text-guided diffusion models in generating molecules with specific properties, highlighting their potential in olfactory research [31] [32]. Further research by Lee, et al. [33] and Sisson, et al. [34] have demonstrated feasibility of other generative models in producing viable chemical compounds according to their desired odours or chemical bonds.
@@
The validation of generated molecules necessitates empirical methods to confirm their olfactory properties. Advanced olfactory sensors offer a means to detect and analyze volatile compounds, providing data on their scent profiles. By comparing the sensor readings of generated molecules with the intended odour descriptors, researchers can assess the accuracy and relevance of the diffusion model’s outputs. In reality, this empirical validation is very difficult for reasons denoted in the Limitations section.
III Related Work
The integration of olfaction into robotic systems has garnered significant attention, particularly in the domain of robotic odour source localization (OSL). Traditional OSL approaches often rely solely on olfactory cues, which can be unreliable in complex environments due to factors like turbulent airflow and the presence of multiple odour sources. To address these challenges, recent studies have explored the fusion of multiple sensory modalities [35] [36].
For instance, a study introduced a fusion navigation algorithm that combines both vision and olfaction-based techniques. This hybrid approach addresses challenges such as turbulent airflow, which disrupts olfaction sensing, and physical obstacles inside the search area, which may impede vision detection. The methodology includes a custom-trained deep-learning model for visual target detection and a moth-inspired algorithm for olfaction-based navigation. Experimental results demonstrate that this vision and olfaction fusion algorithm significantly outperforms vision-only and olfaction-only methods, reducing average search time by 30% and 54%, respectively [35] [36].
In parallel, advancements in machine learning have facilitated the generation of novel molecules based on textual descriptions. Text-guided diffusion models, such as TextSMOG, have been developed to generate small molecules conditioned on textual inputs. These models enhance both stability and diversity in molecule generation, capturing and utilizing information from textual descriptions effectively [31].
Furthermore, the application of diffusion models in molecular design has been extensively reviewed, highlighting their potential in generating molecules with desired properties. These models operate by simulating the gradual degradation of a data distribution and learning its reverse process to generate new samples. Their success in visual domains has inspired researchers to explore their potential for molecular generative tasks, making them central to molecular design [32]. Work from Lee, et al. in [33] show how graph neural networks can be used to construct principal odour maps for deducing probable aromas for a given molecule. Our work here with graph neural networks strives to align with many of their assumptions, but leverages diffusion to overcome uncertainties associated with changing perceivable aromas among different concentrations of the same compound as exemplified by Longin, et al. [37].
@@ III Related Work
Despite these advancements, challenges persist in accurately associating odours with their correct sources in complex environments. Our proposed methodology builds upon these existing works by integrating vision-language models to extract odour descriptors from images and employing diffusion models to generate corresponding molecular structures. This approach aims to enhance the robot’s ability to disambiguate odour sources, improving navigation and decision-making in environments where olfactory cues are essential.
IV Methodology
IV-A Diffusion Model Selection
@@
Traditional methods for exploring chemical similarity, such as nearest neighbors using molecular fingerprints, are effective for retrieving known molecules from existing datasets. These methods are particularly well-suited for tasks that involve classification, clustering, or similarity-based search. However, they fall short when the goal is to generate entirely new molecules that possess specific olfactory properties — especially combinations of scent descriptors not commonly observed in known databases.
@@
In contrast, diffusion models are generative frameworks capable of sampling novel molecular graphs from a learned data distribution. When trained on molecular structures annotated with multi-label scent descriptors, diffusion models can conditionally generate new candidate molecules that reflect specified olfactory profiles (e.g., floral, musky, woody). This goes beyond retrieval — it enables creation, offering several key advantages:
@@
Unlike k-nearest neighbor methods, diffusion models do not rely on existing database entries. They have generative capability and can synthesize new, chemically valid molecular structures that conform to desired scent characteristics. Diffusion models also support multi-label conditioning, allowing fine control over the generated output. One can, for instance, request molecules that are simultaneously fruity, earthy, and sweet — even if no molecule with that combination exists in the training data. By modeling a distribution over molecular structures, diffusion models allow for sampling of diverse outputs that satisfy the same olfactory constraints, improving the breadth of chemical space explored during generation. The denoising process intrinsic to diffusion models implicitly learns smooth transitions in molecular structure and scent characteristics, enabling interpolation and optimization across scent manifolds. The diffusion framework can naturally be extended to incorporate additional modalities such as synthetic accessibility, toxicity, or volatility, making it suitable for multi-objective molecule design.
@@
IV-B Graph Neural Networks
@@
V Model Architecture
V-A Diffusion Framework
Our generative framework is built on a conditional denoising diffusion model, enhanced by an EGNN architecture. This design is tailored to generate novel molecular structures in accordance with specific olfactory descriptors. Unlike traditional approaches that rely solely on molecular fingerprints or 2D-graph representations, our model incorporates 3D-molecular geometry and jointly learns both atomic identities and bond structures through time-reversible denoising steps.

### v4

# Diffusion Graph Neural Networks and Dataset for Robust Olfactory Navigation in Hazard Robotics

@@ Abstract
Navigation by scent is a capability in robotic systems that is rising in demand. However, current methods often suffer from ambiguities, particularly when robots misattribute odours to incorrect objects due to limitations in olfactory datasets and sensor resolutions. To address challenges in olfactory navigation, we introduce a multimodal olfaction dataset along with a novel machine learning method using diffusion-based molecular generation that can be used by itself or with automated olfactory dataset construction pipelines. This generative process of our diffusion model expands the chemical space beyond the limitations of both current olfactory datasets and training methods, enabling the identification of potential odourant molecules not previously documented. The generated molecules can then be more accurately validated using advanced olfactory sensors, enabling them to detect more compounds and inform better hardware design. By integrating visual analysis, language processing, and molecular generation, our framework enhances the ability of olfaction-vision models on robots to accurately associate odours with their correct sources, thereby improving navigation and decision-making through better sensor selection for a target compound in critical applications such as explosives detection, narcotics screening, and search and rescue. Our methodology represents a foundational advancement in the field of artificial olfaction, offering a scalable solution to challenges posed by limited olfactory data and sensor ambiguities [^1] Code, models, and data are made available to the community at:
https://huggingface.co/datasets/kordelfrance/olfaction-vision-language-dataset
@@ I Introduction
The human sense of smell is a complex and nuanced sensory modality, capable of distinguishing an extensive array of odourants. In recent years, artificial olfactory systems have been developed to detect sources of explosives, illegal drugs, and even perform forensic studies. These systems, such as electronic noses, utilize sensor arrays and pattern recognition algorithms to detect and classify compounds [1, 2, 3, 4].
Figure 1: An example of how our dataset (A-D) and diffusion equivariant graph neural network (DEGNN) model (E-F) helps in olfactory robotics. (A) Images are taken from the COCO dataset and aroma descriptors are taken from the GoodScents and Leffingwell Datasets. (B) The VLM is instructed to identify the objects (from the COCO classes) and the aromas in the images. (C) Labels of identified objects are returned with predicted aromas associated with each label. (D) Aroma-object pairings are constructed. (E) The pairings are input into the proposed diffusion model to generate all near-neighboring molecules. (F) All molecules are correlated to the originating objects and stored in a list. (G) Engineer evaluates which objects and target compounds he/she wants to identify from the list and uses this information to down-select the number of olfaction sensors needed. (H) These sensors are integrated into the olfaction device of the robot. (I) Robot is ready for deployment.
Despite sensor advancements, the development of machine learning models for olfactory perception faces significant challenges, primarily due to the scarcity of comprehensive olfactory datasets and the difficulty in acquiring ground truth labels for the aromas associated with specific objects. Humans describe chemical odourants with lingual descriptors such as fruity and floral, but sensors interpret odourants by their direct chemical reactions, leading to a disparity in data training methods for olfactory sensors (e.g. how does one describe to an olfactory sensor the ”scent” of explosives?). Datasets like those from Leffingwell [5] and GoodScents [6] have been instrumental in providing foundational results. However, due to the complexity in obtaining human-labeled odour descriptors over compounds [7], these datasets are limited in scope, lacking complete coverage of the vast chemical space associated with olfactory stimuli [8] and the necessary human controls to ensure objective measurements such as compensation for genetic variation in olfactory mechanisms [9], environmental context [10], health status [11, 12], age [13], and even potential olfaction system damage [14].
To address this data paucity, we propose a vision-olfaction-language dataset along with an innovative diffusion-based equivariant graph neural network (DEGNN) that allows one to minimize uncertainty whilst constructing olfactory datasets and selecting the appropriate sensors for an olfaction-equipped robot. Our motivation is driven by the desire to teach robots to navigate by scent and triangulate the source of explosives, building on methods from [15, 16, 17, 18]. Our research provides four primary contributions:
@@
    Our model aids in the selection of olfactory sensors for a desired target compound or aroma.
@@
    In result of (1) and (2), our model reduces uncertainty due to the lack of grounding in olfactory robotics tasks.
@@
    In result of (3), our methods directly increase accuracy and hardware efficiency in scent-based navigation.
@@
II Background & Related Work
@@
A significant hurdle in developing machine learning models for olfaction is the scarcity of extensive and high-quality datasets. Existing resources, such as the Leffingwell PMP 2001 [5] and GoodScents [6] datasets, provide valuable information on various chemical compounds and their associated odour descriptors. However, these datasets are limited in the number of compounds they cover and do not encompass the full spectrum of olfactory stimuli. In addition, these datasets were constructed by human evaluators without all of the necessary controls needed to qualify the performance of their olfactory mechanisms leading to uncertainty in their grounding [7, 8]. This is not a fault of the dataset authors, but a known problem in olfaction due to the limited range of human olfactory mechanisms [9, 10, 11, 12, 14]. This limitation hampers the training of robust ML models capable of generalizing across diverse olfactory inputs, especially when many olfactory sensors exhibit better sensitivity than humans [19].
Figure 2: Overview of the olfactory-guided diffusion model architecture. Atom features and 3D coordinates evolve over time via EGNN layers, while bond types are predicted in parallel. The model is conditioned on the olfactory descriptor embeddings.
@@
To address the data scarcity and grounding issues, researchers have explored the use of vision-language models (VLMs) to generate odour descriptors from images. VLMs, trained on large-scale image-text pairs, can infer contextual information from visual inputs and generate corresponding textual descriptions [20] [21]. By applying VLMs to images, it is possible to extract descriptive terms of objects within that may correlate to their olfactory characteristics. However, the reliability of these descriptors is contingent on the VLM’s exposure to olfactory related data during training, the degree of which is unknown [19]. Consequently, the generated descriptors may lack specificity or accuracy in representing odours.
@@
While efforts to construct olfaction-vision-language models (OVLMs) exist [22], their underlying training data and training processes are not well standardized. By analyzing visual content in a standard computer vision dataset like, for example, the COCO dataset [23], VLMs can infer potential olfactory characteristics associated with depicted scenes or objects. For example, if a motorcycle is identified in an image, the VLM will reason that carbon monoxide is likely a present molecule and will list the aromas that describe carbon monoxide given the appropriate prompt. These inferred descriptors serve as a bridge to map visual information to molecules, facilitating the identification of molecules likely present in a given image.
Thus, the reliance on VLMs introduces its own set of uncertainties, particularly regarding the accuracy and completeness of the generated odour descriptors which further contributes to the grounding problem. To mitigate these concerns, we integrate a diffusion model trained on existing olfactory datasets and designed to generate novel molecular structures that correspond to the inferred odour descriptors, effectively expanding the chemical space beyond the limitations of current datasets. By doing so, we contribute to the oolfactory grounding problem by capturing a broader spectrum of odourants and molecules - including those not contained in the training data.
@@
Our motivation in doing so is to obtain more reliable and accurate scent-based navigation in robots. Sensors that rapidly sample the air only screen for a select set of molecules. If a robot is given the task to navigate toward a specific odour, it has to be equipped with olfactory sensors that can target the specific molecules attributed to that odour. By understanding the target molecule’s near neighbors on the odour continuum, one can ensure to integrate all sensors that address these molecules as well. As a simplified example, imagine that an unmanned aerial vehicle (UAV) is instructed to find the source of nitric oxide ($NO$), sometimes an indicator for the presence of explosives. $NO$ can, in some environments, quickly oxidize to $NO_{2}$. If the robot takes substantial time to navigate to the source of $NO$, it could be analyzing the wrong compound by missing the detection window, and at some point, may want to transition to detecting $NO_{2}$. The olfactory sensors (for example, metal oxide sensors) required to detect both $NO$ and $NO_{2}$ are different, and it would be prudent for the engineer to query which near neighbor compounds could be relevant for detection so the UAV could be equipped with such sensors. This issue is exemplified in tasks from [15] [16] [17] [18] [24]. We automate such queries in our dataset construction, such that a separate machine learning model is queried for the nearest neighbors of any compound identified by the VLM. This not only informs better sensor selection, but also partially reconciles erroneous compounds produced by the VLM.
@@
In machine olfaction, the Shape Theory of Olfaction suggests that molecules that exhibit similar structural traits ”smell” similarly [25] [26] [27] . Although there is some support against this theory [28] [29] [30] [31], our proposed method assumes this theory is true, and we construct a model that builds off its principles. For example, if a VLM indicates that the odour descriptors ”fruity” and ”floral” are associated with an object in an image, one can use the GoodScent or Leffingwell datasets to look up particular molecules that exude such aromas. However, since these datasets are not comprehensive of all possible molecules, one can use the proposed model to generate ”near neighbors” of molecules that are structurally similar to those found in the existing datasets.
@@
To bridge the gap between odour descriptors and chemical compounds, diffusion models have emerged as a promising solution. Diffusion models are generative frameworks that learn to produce data samples by iteratively denoising random noise, guided by learned data distributions. In the context of molecular generation, diffusion models can be conditioned on textual inputs, such as odour descriptors, to generate novel molecular structures that correspond to the specified olfactory characteristics. This approach enables the exploration of chemical spaces beyond existing datasets, facilitating the discovery of new compounds with desired scent profiles. Recent studies have demonstrated the efficacy of text-guided diffusion models in generating molecules with specific properties, highlighting their potential in olfactory research [32] [33]. Further research by Lee, et al. [7] and Sisson, et al. [34] have demonstrated feasibility of other generative models in producing viable molecules according to their desired odours or chemical bonds.
@@
The validation of generated molecules necessitates empirical methods to confirm their olfactory properties. Advanced olfactory sensors offer a means to detect and analyze volatile compounds, providing data on their scent profiles. By comparing the sensor readings of generated molecules with the intended odour descriptors, researchers can assess the accuracy and relevance of the diffusion model’s outputs. In reality, this empirical validation is very difficult for reasons denoted in the Limitations section. Work from Lee, et al. in [7] show how graph neural networks (GNNs) can be used to construct principal odour maps for deducing probable aromas for a given molecule. Our work here with GNNs strives to align with many of their assumptions, but leverages diffusion to overcome uncertainties associated with changing perceivable aromas among different concentrations of the same compound as exemplified by Longin, et al. [35].
@@
Despite these advancements, challenges persist in accurately associating odours with their correct sources in complex environments. Our proposed methodology builds upon these existing works by integrating vision-language models to extract odour descriptors from images and employing diffusion models to generate corresponding molecular structures. We aim to enhance the robot’s ability to disambiguate odour sources, improving navigation and decision-making in environments where olfactory cues are essential.
III Methodology
III-A Diffusion Model Selection
@@
Traditional methods for exploring chemical similarity, such as nearest neighbors using molecular fingerprints, are effective for retrieving known molecules from existing datasets. These methods are particularly well-suited for tasks that involve classification, clustering, or similarity-based search. However, they fall short when generating entirely new molecules that possess specific olfactory properties — especially combinations of scent descriptors not observed in known databases.
@@
In contrast, diffusion models are generative frameworks capable of sampling novel molecular graphs from a learned distribution. When trained on molecular structures annotated with multi-label scent descriptors, diffusion models can conditionally generate new candidate molecules that reflect specified olfactory profiles (e.g., floral, musky, woody). This goes beyond simple retrieval — it enables creation and generation and thereby offers several key advantages:
@@
Unlike k-nearest neighbor methods, diffusion models do not rely on existing database entries. They have generative capability and can synthesize new, chemically valid molecular structures that conform to desired scent characteristics. Diffusion models also support multi-label conditioning, allowing fine control over the generated output. One can, for instance, request molecules that are simultaneously fruity, earthy, and sweet — even if no molecule with that combination exists in the training data. By modeling a distribution over molecular structures, diffusion models allow for sampling of diverse outputs that satisfy the same olfactory constraints, improving the breadth of chemical space explored during generation. The denoising process intrinsic to diffusion models implicitly learns smooth transitions in molecular structure and scent characteristics, enabling interpolation and optimization across scent manifolds. The diffusion framework can naturally be extended to incorporate additional modalities such as synthetic accessibility, toxicity, volatility, or gas diffusive characteristics making it suitable for multi-objective molecule design.
@@
III-B Graph Neural Networks
@@
IV Model Architecture
IV-A Diffusion Framework
Our generative framework is built on a conditional denoising diffusion model, enhanced by an EGNN architecture which we denote as Diffusion EGNN, or DEGNN for brevity. This design is tailored to generate novel molecular structures in accordance with specific olfactory descriptors. Unlike traditional approaches that rely solely on molecular fingerprints or 2D-graph representations, our model incorporates 3D-molecular geometry and jointly learns both atomic identities and bond structures through time-reversible denoising steps.